Find the general/particular solution of the following Differential Equations


(Non-Exact D.E)


(t² + 1) dI - (t-It) dt=0


1
Expert's answer
2022-01-31T16:45:10-0500

"\frac{t}{t^2+1}dt=\frac{1}{1-l}dl"


"\frac{ln(t^2+1)}{2}=lnc_1-ln(l-1)"


"(t^2+1)(l-1)^2=c"